OpenStudy (anonymous):

explain what is wrong with the following solution for the equation sin2theta=a in the interval {0,2pi} sin 2 theta =1 sin theta = 1/2 theta = pi/6 or theta = 5pi/6 helllpppp em plz

OpenStudy (loser66):

\(sin(2\theta) =1\rightarrow sin \theta =\dfrac{1}{2}\) is wrong

OpenStudy (loser66):

it should be sin (\(2\theta)\) = 1 \(\rightarrow 2\theta = arcsin (1)= \dfrac{\pi}{2}\) so, \(\theta = \dfrac{\pi}{4}\)

OpenStudy (loser66):

other solution is 3pi/4

OpenStudy (anonymous):

thanks can u help me with one more plz

OpenStudy (anonymous):